OpenStudy (anonymous):

what is the last number in the series number 7,21,63,189?

OpenStudy (anonymous):

189*3

OpenStudy (anonymous):

Well as you can see from the other numbers it seems it is multiply be 3 For example 7 * 3 = 21, 21 * 3 = 63, 63 * 3 = 189 So if you are trying to find the next number just multiply 189 * 3 = 567

OpenStudy (anonymous):

567

OpenStudy (agent0smith):

"what is the last number in the series number" "last number" isn't very descriptive here... you mean the next number?

OpenStudy (anonymous):

tnx..